Actions
Physics Shift Report¶
Start Date: 7 June
End Date: 11 June
Start Run: 6628
End Run: 6896
Physics shifter: Paolo Franchini
Summary¶
Purpose of the shifts was- taking data to figure out tracker timing
- taking data to test Run Control
- monitor and fix the online reconstruction
- TOF Calibration run
Run documentation¶
- Run plans: http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlans
- TOF calibration: http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20150611#TOF-Calibration-run
Elog¶
Nothing to comment.
Run Condition Summary Spreadsheet/CDB¶
If it was used during the shift, has still to be uploaded here: http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunSummarySpreadsheets
Analysis¶
Plots almost completely empty until the 11/6 night shift (http://micewww.pp.rl.ac.uk/analysis/physics_shifts/6800.html) where we finally saw slabs on TOF0, TOF1 and TOF2 and also timing plots for TOF0, TOF1 and TOF0->TOF1.
Updated by Franchini, Paolo over 8 years ago ยท 13 revisions